OpenStudy (anonymous):

tell me how to solve this [indefinite integral]sqrt(a+bsin^2(theta)) d(theta), a and b are here real constants.

myininaya (myininaya):

\[\int\limits_{}^{}\sqrt{a+bsin^2(\theta)} d \theta\]

myininaya (myininaya):

\[\int\limits_{}^{}\sqrt{a}\sqrt{1+\frac{b}{a}\sin^2(\theta)} d \theta\]

myininaya (myininaya):

\[\sqrt{a}\int\limits_{}^{}\sqrt{1+\frac{b}{a} \sin^2(\theta)} d \theta \]

myininaya (myininaya):

\[\sqrt{a}\int\limits_{}^{}\sqrt{1+(\sqrt{\frac{a}{b}} \sin (\theta))^2} d \theta\]

myininaya (myininaya):

thinking...

OpenStudy (anonymous):

I actually stuck in that step no subsitution or parts work here

myininaya (myininaya):

maybe this might work let \[\tan(x)=\sqrt{\frac{b}{a}}\sin(\theta)\]

myininaya (myininaya):

\[\sec^2(x) dx=\sqrt{\frac{b}{a}} \cos(\theta) d \theta\]

myininaya (myininaya):

nope

OpenStudy (anonymous):

i don't think so I appliled subsitution like that also did'nt work

OpenStudy (anonymous):

please see my one more question. I think you can help me.

myininaya (myininaya):

what?

myininaya (myininaya):

i don't see another question

OpenStudy (anonymous):

I posted but no body try to answer that question you can see in my profile

OpenStudy (anonymous):

I want to know how it's solved

OpenStudy (zarkon):

it's an elliptic integral ... good luck

OpenStudy (anonymous):

so how it's solved